OpenStudy (anonymous):

Arthur is trying to find the equation of a perpendicular to y = 1 over 4x + 4 in slope-intercept form that passes through the point (−2, 6). Which of the following equations will he use?

OpenStudy (anonymous):

What are the following equations @imclassylikethat

OpenStudy (anonymous):

y − 6 = 1 over 4(x − (−2)) y − (−2) = 1 over 4(x − 6) y − 6 = −4(x − (−2)) y − (−2) = −4(x − 6)

OpenStudy (anonymous):

Ok cool thanks

OpenStudy (welshfella):

Hint if the slope of a line is say 2 then the slope of a line perpendicular to it is -1/2

OpenStudy (welshfella):

the slope of the line in the question is 1/4 so what about the slope of its perpendicular?

OpenStudy (anonymous):

-4/1

OpenStudy (anonymous):

@welshfella It will either be y-6=-4(x-(-2)) or y-(-2)=-4(x-6)

OpenStudy (welshfella):

you are right about the slope - its -4 to get the full equation use the standard form y - y1 = m(x - x1) where m = slope = -4 and x1 anf y1 are the coordinates of the given point that is x1 = -2 and y1 = 6

OpenStudy (anonymous):

So it's y-(-2)=-4(x-6) correct?

OpenStudy (welshfella):

no y1 = 6

OpenStudy (anonymous):

Okay. thank you. Can you help me with some other problems?

OpenStudy (welshfella):

I only have about 10 minutes before i have to go so i can help until then

OpenStudy (anonymous):

I need the perimeter of this shape

OpenStudy (welshfella):

some lengths can just be read off directly - those which are vertical or horizontal

OpenStudy (welshfella):

but you;ll have to use the pythagoras theorem to find the lengths of of the sloping lines

OpenStudy (welshfella):

for example the line SR